1) Which of the following processes would be expected to have the largest increase in ?Srxn per mole of benzene?

a) C6H6(l) + Cl2(g) ?  C6H5Cl(l) + HCl(g)

b) C6H6(l) + 15/2 O2(g) ? 6CO2(g) + 3H2O(g)

c) C6H6(s) ? C6H6(g)

d) 6H6(s)  ? C6H6(l)

e) C6H6(l) + 9/2 O2(g) ? 6CO(g) + 3H2O(g)

2) The heat of fusion of acetic acid is 11.5 kJ/mol. Its melting point is 16.6 oC. The change in entropy for the melting of acetic acid in J/(mol.K) is:

a)11.5
b)39.7
c)694
d)-11.5
e)15.1

3) ?S is positive for the reaction __________.

a)CO2 (g) ? CO2 (s)
b) 2Hg (l) + O2 (g) ? 2HgO (s)
c) BaF2 (s) ? Ba2+ (aq) + 2F- (aq)
d)2NO2 (g) ? N2O4 (g)
e)2H2 (g) + O2 (g) ? 2H2O (g)
Answer #1

Entropy determines the randomness of the particle.

So gases will have more Entropy than liquids and liquid will have more entrEnt than solid.

1. C6H6 + 15/2 O2 = 6CO2(g) + 3 H2O (g)

Will have more change in entropy . It has the maxomax no. of gaseous molecules.

CO2 will have more Entropy than CO because it has more no of atoms in a molecule

2. Entropy = ?H / T = 11.5*10^3 / (16.6+273) = 39.709 J/K

3. Option C is correct

BaF2 is solid whereas Ba2+(aq) and F-(aq)

If you satisfied with the solution please like it thankxx